# Physical-Layer Machine Learning with Multimode Interferometric Photon Counting

### v1

@@ IV Results and Comparison
$A=\absolutevalue{\vec{w}^{T}\vec{w}^{\star}}^{2}.$ (IV.1)
This metric provides a direct measure of how closely the variational circuit approaches the true solution. In contrast to the loss function, which can fluctuate due to finite-shot noise, the accuracy provides a better indication of convergence to the ground truth. As we now elaborate, photon counting consistently yields higher accuracy and more robust training than homodyne detection for both PCA and CCA tasks. To begin, we focus on a fixed value of signal strength $\sigma_{c}=0.02$ and number of modes $M=21$ and present the training history of the accuracy versus time steps $t$ in Fig. 2. In subplots (a)(c), we see a substantial increase of the accuracy as training proceeds in the case of photon counting, despite fluctuations coming from noises in loss function estimation. In contrast, (b)(d) shows a poor convergence of the training for homodyne detection.
@@
We next study the effect of scaling the number of modes $M$, while holding the total signal strength fixed at $\sqrt{M}\sigma_{c}=\text{const}$. In principle, this normalization preserves the signal-to-noise scaling of the problem, since we consider completely correlated noise (i.e., rank 1 covariance, $\bm{V}$, with eigenvalue $M\sigma_{c}^{2}$). However, in practice, the optimization landscape becomes more complex as $M$ increases due to an increase in the number of parameters, and the training process may converge to suboptimal local minima. As shown in Fig. 4, the resulting accuracy decreases with increasing $M$, reflecting this growing optimization complexity.
Figure 4: Accuracy as a function of the number of modes $M$ for (a) PCA and (b) CCA. Each data point corresponds to a separate training run with distinct initialization and signal sampling. We fix $\sqrt{M}\sigma_{c}=0.2$ to maintain constant total signal strength. Dashed lines indicate the expected random guess limit $1/M$.
@@
V Further enhancement from squeezing
@@
The distributed squeezed-vacua assisted variational quantum processor has a similar flavor to the original SLAEN recipe [8] but, crucially, we augment the detection stage to include anti-squeezing operations and photon counting measurements (see also Refs. [25, 26] for its relevance in single-mode parameter estimation). This modification at the receiver end enhances sensitivity to weak stochastic signals, surpassing the performance achievable with homodyne detection considered in Refs. [8, 9]. The induced advantage of the proposed scheme over the original SLAEN is identical to the advantage of photon counting over homodyne, due to the amplification of signals applying to both current protocol and original SLAEN.
@@ VI Conclusion
The proposed scheme bears conceptual similarity to the SLAEN framework [8, 9], but crucially differs in its detection strategy. Whereas SLAEN employed homodyne measurements with classical post-processing, our approach incorporates programmable linear-optics directly at the measurement stage, followed by photon counting. This tradeoff — replacing classical post-processing with physical-layer variational quantum processing and homodyne detection with photon counting — yields a substantial gain in sensitivity and performance.
@@
The methods developed here are well-suited to emerging application domains in quantum-enhanced estimation and learning. These include distributed quantum sensing networks for fundamental physics applications [29, 30, 31], such as searching for dark matter [16, 17, 32, 33] and stochastic gravitational wave signatures [34, 35].
@@
Our variational estimation approach may also prove useful in the seemingly disparate domain of sub-diffraction incoherent imaging [36, 37], where optimal performance is known to rely on linear optics and photon counting. In this context, learning the first few principle components of a complex multi-emitter scene through variational quantum measurements may provide a practical and scalable path toward dispelling “Rayleigh’s curse” [38].

### v2

@@ IV Results and Comparison
[…]
This metric provides a direct measure of how closely the variational circuit approaches the true solution. In contrast to the loss function, which can fluctuate due to finite-shot noise, the accuracy provides a better indication of convergence to the ground truth. As we now elaborate, photon counting consistently yields higher accuracy and more robust training than homodyne detection for both PCA and CCA tasks. To begin, we focus on a fixed value of signal strength $\sigma_{c}=0.02$ and number of modes $M=21$ and present the training history of the accuracy versus time steps $t$ in Fig. 2. In subplots (a)(c), we see a substantial increase of the accuracy as training proceeds in the case of photon counting, despite fluctuations coming from noises in loss function estimation. In contrast, (b)(d) shows a poor convergence of the training for homodyne detection. Note that even for the cases with photon counting in subplots (a)(c), the convergence towards unity accuracy is ultimately impeded by the measurement noise from finite sampling.
@@
We next study the effect of scaling the number of modes $M$, while holding the total signal strength fixed at $\sqrt{M}\sigma_{c}=\text{const}$. In principle, this normalization preserves the signal-to-noise scaling of the problem, since we consider completely correlated noise (i.e., rank 1 covariance, $\bm{V}$, with eigenvalue $2M\sigma_{c}^{2}$). However, in practice, the optimization landscape becomes more complex as $M$ increases due to an increase in the number of parameters, and the training process may converge to suboptimal local minima. As shown in Fig. 4, the resulting accuracy decreases with increasing $M$, reflecting this growing optimization complexity.
Figure 4: Accuracy as a function of the number of modes $M$ for (a) PCA and (b) CCA. Each data point corresponds to a separate training run with distinct initialization and signal sampling. We fix $\sqrt{2M}\sigma_{c}=0.2$ to maintain constant total signal strength. Dashed lines indicate the expected random guess limit $1/M$.
@@
V Effects of Practical Imperfections
In this section, we analyze departures from the idealized model, including photon-detector inefficiency and dark counts, dephasing in the linear-optical unitary, and non-Gaussian signals.
V.1 Detector Efficiency and Dark Counts
Figure 5: Average training performance of PCA in the presence of dark counts. Curves in different colors correspond to varying levels of effective thermal noise induced by dark counts ($n_{\rm DC}=\Gamma_{\rm DC}\tau$). The number of modes is fixed at $M=21$. Vertical dashed lines indicate the points where the signal-to-noise ratio (SNR) reaches 3% under different thermal noise levels.
Photon detectors exhibit imperfections that significantly degrade estimation performance, particularly in the weak-signal regime. The primary culprits are finite detection efficiency and dark counts — spurious detection events arising from background radiation or electronic instabilities. We assess the impact of such imperfections on PCA (similar analyses hold for CCA), presenting a simplified theoretical model alongside supporting numerical results.
Real photon detectors suffer from two primary imperfections: finite quantum efficiency and dark counts. Finite efficiency, denoted $\eta$, models the probability that an incident photon is successfully detected. Whereas dark counts arise from thermal or electronic noise and can be modeled as background excitations that appear in the absence of the signal.
We mathematically model an imperfect photon detector by a three-stage process: first, an attenuation channel with transmission $\eta$; second, a random displacement $\beta\sim\mathcal{CN}(0,n_{\rm DC})$ representing additive noise from dark counts, where $n_{\rm DC}=\Gamma_{\rm DC}\tau$ is the expected number of dark counts over an integration time $\tau$, $\mathcal{CN}$ is complex normal distribution; and third, ideal photon-number detection.
The input state to the imperfect photon detector is the output of the PCA variational processor, which is a single-mode, incoherently displaced state with displacement $\lambda_{w}\sim\mathcal{N}(0,\vartheta_{\rm PCA})$ along the $P$ quadrature [recall $\vartheta_{\rm PCA}=\vec{w}^{T}\bm{V}\vec{w}$]. After accounting for attenuation and dark counts, the effective state remains an incoherently displaced state, but now with variance $\eta\vartheta_{\rm PCA}+n_{\rm DC}/2$ along the $P$ quadrature and variance $n_{\rm DC}/2$ along the $Q$ quadrature due to dark counts affecting both quadratures equally. The average photon counts measured by the detector is thus $\bar{n}=\eta\vartheta_{\rm PCA}/2+n_{\rm DC}$. The first term originates from the signal and the second from the dark-count noise, which represents the detector noise floor.
In the presence of detector inefficiency and dark counts, the loss observable must be augmented as
$\hat{L}=-\frac{2}{\eta}(\hat{N}_{w}-n_{\rm DC}),$ (14)
so that $\expectationvalue*{L}=-\vartheta_{\rm PCA}$, as before. In other words, the dark counts must be characterized and subtracted off, while the finite quantum efficiency requires rescaling of the estimator. Furthermore, the loss observable now suffers higher noise fluctuations. In the regime where dark counts dominate the total photon number, $\bar{n}\approx n_{\rm DC}$, we find that
$\mathrm{Var}(\hat{L})\approx\frac{4}{\eta^{2}}n_{\rm DC}(n_{\rm DC}+1).$ (15)
[…]
$\frac{n_{\rm DC}(n_{\rm DC}+1)}{\eta^{2}}\ll\frac{1}{8}.$ (16)
For detectors with $\eta\gtrsim 0.9$ and dark-count rates $n_{\rm DC}\ll 1$, this condition is easily satisfied. Note that this analysis pertains to the physical estimation noise. Though how this translates to adaptively learning (e.g., learning the PCA vector $\vec{w}^{\star}$) is non-trivial. We provide quantitative numerical assessments to address this below.
Here we numerically investigate how dark counts affect photon counting performance. For simplicity, we assume perfect detection efficiency ($\eta=1$).
For the numerical analysis, we model the detector statistics as a Poisson process. In the absence of an input signal, the detector receives an average of $n_{\rm DC}$ spurious photons per unit time. Note that the signal (i.e., the output of the quantum processor) is a random coherent state with mean photon number per unit time $\bar{n}_{\rm sig}=\vartheta_{\rm PCA}/2$. So that, in the ideal case of no dark counts, the photon detection statistics follow a Poisson distribution with mean $\bar{n}_{\rm sig}$, i.e., ${\rm Poisson}(\bar{n}_{\rm sig})$. When dark counts are included, this distribution is effectively shifted to ${\rm Poisson}(\bar{n}_{\rm sig}+n_{\rm DC})$.
In Fig. 5, we take PCA as an example to illustrate how performance degrades with the dark-count rate. Note that the average photon number $\bar{n}_{\rm sig}$ at the output is proportional to the raw signal $M\sigma_{\rm c}^{2}$. [^4] Recall that $\bar{n}_{\rm sig}=\vartheta_{\rm PCA}/2=\vec{w}^{T}\bm{V}\vec{w}/2$ and that $\bm{V}=2M\sigma_{c}^{2}\vec{v}\vec{v}^{T}$. When $\bar{n}_{\rm sig}\gg n_{\rm DC}$, the effect of thermal noise is negligible, and the performance approaches that of the ideal detector. In contrast, when $\bar{n}_{\rm sig}\ll n_{\rm DC}$, noise dominates and photon-counting accuracy deteriorates sharply.
We use the dashed line in Fig. 5 to indicate the signal strength at different dark-count levels, where the analytical calculations of the SNR coincide. The ratio between the signal amplitude and the noise amplitude is 3%. The accuracy remains similar for different dark-count levels around 0.1, which marks the approximate lower bound of signal strength required to achieve a quantum advantage. Beyond this threshold, the accuracy improves significantly.
In practice, as long as the dark-count rate is sufficiently low such that $n_{\rm DC}<\bar{n}_{\rm sig}$, photon counting remains effective despite the presence of detector noise.
V.2 Dephasing
Figure 6: Accuracy as a function of dephasing strength for PCA. The dephasing strength is $\sigma_{\varphi}=0.02$. Other parameters are $M=21$ and $\sigma_{c}=0.02$.
[…]
In the case of photon counting, the dephased signal strength for PCA becomes
$2N_{w}$ $=$ $\left|\sum_{l}w_{l}\lambda_{l}\cos\varphi_{l}\right|^{2}+\left|\sum_{l}w_{l}\lambda_{l}\sin\varphi_{l}\right|^{2}$ (17)
$=$ $\sum_{l}\left|w_{l}\lambda_{l}\right|^{2}+\sum_{k\neq l}w_{k}\lambda_{k}w_{l}\lambda_{l}\cos(\varphi_{k}-\varphi_{l})~.$
For homodyne detection, only momentum quadrature is measured. The corresponding signal strength of homodyne detection for PCA after dephasing becomes
$P_{w}$ $=$ $\sum_{l}w_{l}\lambda_{l}\cos\varphi_{l}$
$P_{w}^{2}$ $=$ $\left|\sum_{l}w_{l}\lambda_{l}\cos\varphi_{l}\right|^{2}~.$ (18)
In the limit $\varphi_{l}\ll 1$, the signal strength is reduced by approximately $O(\varphi_{l}^{2})$, indicating that our protocol is not overly sensitive to dephasing. Furthermore, since Eq. (17) exceeds Eq. (18), photon counting yields a stronger signal than homodyne detection.
We assume that iid dephasing in each mode, with each following a Gaussian distribution $\varphi_{l}\sim\mathcal{N}(0,\sigma_{\varphi})$. In Fig. 6, we illustrate the weak-signal case for PCA. The results show that photon counting remains robust against weak dephasing, as indicated by the extended plateau of the curve. A noticeable drop in accuracy occurs only when the dephasing becomes very strong.
V.3 Non-Guassian Signals
In our benchmark problems, the signals are modeled as Gaussian displacements, where the covariance matrix fully characterizes the data. This assumption, however, neglects the possibility of non-Gaussian statistics, which are often relevant in practical sensing scenarios. To illustrate this, we consider a bimodal Gaussian distribution, $\lambda\sim\left(\mathcal{N}(-\sqrt{10M}\sigma_{c},\sqrt{2M}\sigma_{c})+\mathcal{N}(\sqrt{10M}\sigma_{c},\sqrt{2M}\sigma_{c})\right)/2$. As shown in Fig. 7, our results demonstrate that the proposed method continues to perform effectively even in the presence of such non-Gaussian signals.
Figure 7: Accuracy versus non-Gaussian signal strength for PCA. Other parameters are $M=21$ and $\sigma_{c}=0.02$.
VI Further enhancement from squeezing
@@
The distributed squeezed-vacua assisted variational quantum processor has a similar flavor to the original SLAEN recipe [8] but, crucially, we augment the detection stage to include anti-squeezing operations and photon counting measurements (see also Refs. [37, 47] for its relevance in single-mode parameter estimation). This modification at the receiver end enhances sensitivity to weak stochastic signals, surpassing the performance achievable with homodyne detection considered in Refs. [8, 9]. The induced advantage of the proposed scheme over the original SLAEN is identical to the advantage of photon counting over homodyne, due to the amplification of signals applying to both current protocol and original SLAEN. Including variational quantum processing at the front-end is closer in spirit to a full variational quantum toolbox [25].
@@ VII Outlook
The proposed scheme bears conceptual similarity to the SLAEN framework [8, 9], but crucially differs in its detection strategy. Whereas SLAEN employed homodyne measurements with classical post-processing, our approach incorporates programmable linear-optics directly at the measurement stage, followed by photon counting. This tradeoff — replacing classical post-processing with physical-layer variational quantum processing and homodyne detection with photon counting — yields a substantial gain in sensitivity and performance. This approach is also reminiscent of structured receivers developed for coherent-state discrimination [48, 49, 50].
@@
The methods developed here are well-suited to emerging application-specific domains in quantum-enhanced estimation and learning. These include distributed quantum sensing networks for fundamental physics applications [53, 54, 55], such as searching for dark matter [18, 19, 40, 36] and stochastic gravitational wave signatures [56, 57] (see Section II.4.1 for further discussion).
@@
Quantum sensing with a variational network of mechanical oscillators, such as trapped ions [39] or optomechanical systems [14, 40, 15], provides a natural playground for our methods. Stochastic forces or fields with rich spatial distributions can drive the sensor array, and identifying the most informative collective modes is paramount. Assessing correlations between collective modes in the array may be interesting from a physics viewpoint.
Our variational approach may also prove useful in the domain of sub-diffraction incoherent imaging [31, 33, 34, 35], where optimal performance is known to rely on linear optics and photon counting. In this context, learning the first few principle components of a complex multi-emitter scene through variational quantum measurements may provide a practical path towards dispelling “Rayleigh’s curse” [32].
acknowledgements
QZ acknowledges discussions with Zheshen Zhang. This work is supported by Office of Naval Research Grant No. N00014-23-1-2296. QZ also acknowledges support from National Science Foundation Grant No. 2350153 and Defense Advanced Research Projects Agency (DARPA) D24AC00153-02.
Data availability
The codes and data that support the findings of this article are available at [58].